OpenStudy (anonymous):

Fun question time: If a certain type of bacteria doubles every minute, and takes an hour to fill a jar, how long will it take 2 of that bacteria to fill the same jar?

OpenStudy (amistre64):

B{0} = 1; B{n} = 2(B{n-1}) B{1} = 2(1) B{2} = 2(2(1)) B{3} = 2(2(2(1))) = 2^3 B{4} = 2^4 B{60} = 2^60 = full jar :) for starters

OpenStudy (amistre64):

B{0} = 2 ; B{n} = 2B(n-1) B{1} = 2(2) B{2} = 2(2(2)) B{3} = 2(2(2(2))) = 2^4 B{n} = 2^(n+1) 2^60 = 2^(n+1) when n+1 = 60; n=59 minutes perhaps?

OpenStudy (anonymous):

thats correct! That was the first way I solved it as well, from a mathematical standpoint. Then after thinking about it a little more, I came up with a logical argument. Here are the two solutions.

OpenStudy (amistre64):

yep :)

OpenStudy (amistre64):

but as we all know; logic has no place in math :)

OpenStudy (anonymous):

hahaha